OpenStudy (anonymous):

PLEASE HELP WILL GIVE MEDAL AND FAN Find the roots of the polynomial equation. x^2 + x - 72 = 0 9 and -8 -6 and 6 -2 and 16 -9 and 8

OpenStudy (anonymous):

the roots are -9 and 8 because when you factor the equation it gives you (x+9)(x-8)=0. then you set x+9=0, x=-9 and set x-8=0, x=8. therefore, the roots are -9 and 8 :)

OpenStudy (anonymous):

Thank u